OpenStudy (anonymous):

ΔABC is similar to ΔDEF. The ratio of the perimeter of ΔABC to the perimeter of ΔDEF is 1 : 10. If the longest side of ΔDEF measures 40 units, what is the length of the longest side of ΔABC ? 2 units 4 units 16 units 30 units

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

again set up a proportion using the given ratio 1/10 \[\frac{x}{40} = \frac{1}{10}\]

OpenStudy (anonymous):

please just help me> im really just finna give up. i dont even have time to learn the stuff on plato. i just need answers or ima fail:(